Andy is investing $100,000 and wants to accumulate $125,000. He will earn a return of 6% per annum payable half yearly. How many years will it take for Andy to save up to $125,000.

The answer is 4 years

I can't get the answer as 4 how?

Formula is S=P(1+i)^n
Answer #1

PV = 100000
FV = 125000
Rate semi annually = 6%/2 = 3%
Let number of Periods = 2*n =2n
FV = PV*(1+r)2n
125000 = 100000*(1+3%)2n
1.25 = 1.032n
applying log on both sides we get
log(1.25) = 2n*log1.03
n = 1/2*log(1.25)/log(1.03) = 3.77 years hence No of years is close to 4 years.

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well